OpenStudy (anonymous):

In a right triangle, find the length of the side not given b=1, c=√15

OpenStudy (anonymous):

Does it say is the side is the one not adjacent to the right angle? If so, then a = sqrt(b^2 + c^2) = sqrt(1 + 16) = sqrt(16) = 4 sqrt = square root

OpenStudy (anonymous):

it has an a side but it is a triangle

OpenStudy (anonymous):

What I meant is: is c the longer side or one of the two short ones? If you have a figure, it would help.

OpenStudy (anonymous):

i dont have a graph b=1, c=√5

OpenStudy (anonymous):

My answer should be fine

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Two triangles with the following sides:\[\left\{4,\sqrt{15},1\right\} \text{ or } \left\{\sqrt{15},\sqrt{14},1\right\} \]